OpenStudy (anonymous):

Find the value of the discriminant and determine the number of real-number solutions for this quadratic equation. b2-12b+36=0

OpenStudy (anonymous):

Since quadratic equation has form ax^2+bx+c=0 its discriminant is sqrt(b^2-4*a*c), your a=1, b=-12, c=36, the value of discriminant we're looking for is D = sqrt(144-4*36), D = sqrt(0). Therefor, your discriminant is zero and you have only one real solution to this equation which is x = 6 :)